OpenStudy (anonymous):

A glass case is in the shape of a rectangular prism. The volume of the case is 4/125 cubic foot. How many cubic blocks with a side length of 1/5 foot would be required to find the volume of the glass case? Please help I will give medals. A. 2 B. 3 C. 4 D. 5

OpenStudy (anonymous):

start by finding the volume of the cubic blocks. \[V = s^3\] where s = 1/5

OpenStudy (anonymous):

then divide the given volume of the case by the one you found to get the answer

OpenStudy (anonymous):

i got 1/25

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

did you do this?